Links: Minutes, PSB home page, PSB Supervisor ROTA
PRESENT: M. Benedikt, C. Carli, G. Cyvoct, D. Küchler, J-L. Sanchez-Alvarez, R. Scrivens, K. Schindl, M. Vretenar.
Supervisor Linac: R. Scrivens à M. Vretenar
Supervisor PSB: G. Cyvoct à M. Benedikt
Thursday: 20:35 - Valve LI.VVS10 shut after a flashover – would not respond to control at first.
Friday: 05:15 - Power cut. 10:30 - Water back. 10:50 - Linac 2 back, manual reset of 68 tank quadrupole power supplies, trip switch on LTB DC quads.
Sunday: 02:37 - Network hub lost (all DSC could not communicate, causes the watchdog to stop the beam) then a reboot of DSCs meant they were all down. Network specialist came, then the Control network specialist (who doesn't exist anymore). Luckily, Georges- Henri came on shift and remembered this problem (possibly from the PSB). 06:30 - Network and DSCs ok, but Linac 2 would not work, 07:30 - Beam back - Vistar was "locked" so could not see any beam. Also RFQ not tuned so no transmission.
Sunday 16:37 - Network hub controller failed a second time. Reset by operator
Sunday 20:00 - After flashover the CAN bus or a DAC on the HT platform locked up and caused erratic gas flow. Another intervention required.
Monday: 20:11 – Network problem again.
Friday: 14:00 Ions available after power cut.
Friday: 16:00 Transformer head amplifier changed on LTB.MTR55
Saturday: 11:00 - MicroWave generator down due to over heating, and air-conditioning in building 351 off (30oC), repaired by TCR. 17:30 - Linac 3 back to 25oC – Start source switch on, at 19:30 - ECR back to a reasonable intensity.
Sunday ??:?? - Linac 3, water interlock of MicroWave generator, but the flow does not drop. 13:30 - ECR back to reasonable intensity.
Monday: 15:20 - Change of stripper in the Linac gives 25% more intensity in the PSB. There is suspicion of crystallization of the carbon, changing of thickness and warping – which influence the beam parameters. We must verify that there are enough foils to last the run!
To stop the false interlock signals, we propose to swap between contacts on the Microwave generator flow meter on Thursday (time to be decided with M O’Neil is back on Thursday).
The lead sample should be changed before the Ion run, propose Sunday 22 Sept.
Wednesday, 17:30: ISOLDE requested 1 GeV beam to the HRS target without forewarning. Since this was the first request of a 1 GeV HRS beam during the 2002 run, the PSB operations team had no possibility for setting up of the (non ppm) line before. As a consequence the horizontal bending magnet BTY.BHZ308 suffered from beam losses that were detected by TIS on Monday. The ISOLDE team should remember to give an early notice and at least 2 hours of DEDICATED time (during NORMAL WORKING HOURS) to the PSB operations team for setting up of the non-ppm transfer lines before launching non standard beam requests when no up-to-date archives exist.
Friday, 05:15: Power cut. After having the cooling water available (10:30) it took about 1h30 to get the PSB restarted. Beam was ok from 12:00. The main power supply BR.MPS had to be restarted by the PIPO, the power supply of the bending magnet BT.BHZ10 was switched to the spare unit.
On Monday morning we found that the spare BT.BHZ10 supply gives a ~2A lower current than the normal supply for the same ccv value. This caused oscillations of the trajectories in the BTP line. During the Wednesday MD the PO-team will switch back to the normal unit.
Pb ions: On Wednesday about 4 10^10 charges were injected into the PSB. The life time measured was 80 - 100 ms for the 4 rings. On Monday beam was accelerated on all rings, about 1 10^10 charges at ejection. Due to the power cut, the sublimation pumps were off since Friday, they will be switched back on during the Wednesday MD.
Minutes by Michael Benedikt.